# Dynamic Classification: Leveraging Self-Supervised Classification to Enhance Prediction Performance

### v1

@@ 6 Summary and Next Steps
This paper introduces a dynamic classification algorithm (DCA) designed to address the limitations of existing prediction models. By leveraging self-supervised learning based on the distribution of sample data in the training set, the algorithm effectively narrows down the range of data within each subset, leading to improved prediction accuracy. The algorithm consists of three key components: a regular part, a classification part, and a prediction part. Each categorized subclass, containing both predicted and training data, allows the algorithm to predict smaller ranges independently, thus enhancing overall prediction results. As a consequence, DCA is highly sensitive to classification errors, and when these errors are minimized, the algorithm demonstrates excellent performance.
In the case of non-public datasets, particularly factory production data, the DCA achieves a classification error rate of less than 1%, significantly enhancing prediction accuracy. For public datasets, while the classification errors are slightly higher, the performance of DCA remains comparable to other advanced models, including Random Forest and XGBoost. Notably, in the Estimation Obesity dataset, where the classification error rate is only 1.4%, DCA delivers exceptional results, confirming its ability to utilize the first part of supervised classification to effectively support data filtering. This, in turn, allows the algorithm to outperform existing regression models.
Despite its promising results, the DCA algorithm still requires optimization, particularly in the classification phase. Future research will focus on three primary optimization areas:
@@
    Automated Parameter Tuning: Developing an automated system to dynamically adjust the number of classifications and sensitivity based on the data distribution, ensuring more precise partitioning for improved classification performance.
@@
    Classification Model Dependency: Enhancing the algorithm’s robustness by reducing its dependence on the classification model’s effectiveness. A stronger classification model will lead to better results, particularly when the feature correlation is weak.
@@
    Avoiding Local Optima: Introducing mechanisms to prevent the algorithm from falling into local optima, ensuring that the model converges to the most optimal solution.
These optimizations will further improve the robustness and applicability of the DCA algorithm, enabling it to handle a wider range of data characteristics and real-world scenarios, and provide even more reliable and accurate predictions.

### v2

@@ 5.2 Open Source Data
In our previous research, we demonstrated that introducing a new model to achieve zero false negatives and minimal false positives was an unwise choice, both in terms of model performance and training set requirements, which led us further from our objective. Closed-source data typically originates from manufacturers with stable production processes and strong management, which results in higher data purity and consistency, generally yielding higher-quality data. Such data are often well-standardized and consistent, making them more suitable for direct analysis and modeling, typically leading to reliable outcomes. In contrast, open-source datasets are generally associated with greater uncertainty because they may originate from various production or testing environments, with the quality of data potentially varying significantly between sources. Consequently, open-source datasets tend to be more complex and challenging. Therefore, this part of the research focuses on analyzing the impact of feature exclusion versus inclusion using the DCA model on prediction performance.
The DCA performed admirably across all four datasets, achieving top-tier results, particularly when the classification error was below 5%. However, because the DC algorithm relies on a classification method for subclass prediction, its performance is impacted when the feature correlation is weak, leading to higher classification errors. In such cases, even after some data are excluded, the results only approach the performance of Random Forest and XGBoost. When analyzing the Estimation Obesity dataset, which benefits from stronger feature support, the classification error was very low. This indicates that the samples in each subset are nearly always accurate, thereby allowing the $OptimalSegmentationList$ to effectively support the culling function. This enables data reorganization without requiring additional models, thereby significantly improving prediction results. Conversely, datasets such as appliances energy prediction and student performance prediction, which have higher classification errors, lack strong feature support, making it difficult to surpass the performance of Random Forest and XGBoost. Nonetheless, the DCA shows great potential, particularly when classification errors are low, allowing it to filter out low-accuracy data through hyperparameter design and the culling function, thereby enhancing the final prediction results.
@@ 6 Summary and Next Steps
This study introduces the DCA that designed to address the limitations of existing prediction models. By leveraging self-supervised learning based on the distribution of sample data in the training set, the algorithm effectively narrows the range of data within each subset, leading to improved prediction accuracy. The algorithm consists of three key components: regularity, classification, prediction. Each categorized subclass, containing both predicted and training data, allows the algorithm to independently predict smaller ranges, thereby enhancing the overall prediction results. Consequently, DCA is highly sensitive to classification errors, and when these errors are minimized, the algorithm demonstrates an excellent performance. In the case of non-public datasets, particularly factory production data, DCA achieved a classification error rate of less than 1%, significantly enhancing the prediction accuracy. For public datasets, although the classification errors were slightly higher, the performance of the DCA remained comparable to that of other advanced models, including Random Forest and XGBoost. Notably, in the Estimation Obesity dataset, where the classification error rate was only 1.4%, DCA delivered exceptional results, confirming its ability to utilize the self-supervised information of DCA to effectively support data filtering. In turn, this allowed the algorithm to outperform existing solutions.
Despite its promising results, the DCA algorithm still requires optimization, particularly in the classification phase. Future resea rch will focus on these three primary optimization areas:
@@
    Automatic Parameter Adjustment: Develop an automated system capable of dynamically adjusting the number and sensitivity of classifications according to the data distribution. This adjustment aims to achieve more accurate segmentation and improve the classification performance. In addition, explore the possibility of designing a mechanism that enables the DCA to flexibly select the prediction model for each subset within the Redundant Prediction section. The integrated hyperparameter framework is also theoretically feasible. Currently, a fixed prediction model is manually selected, which may limit the prediction accuracy. By enabling flexible model selection, better prediction results can potentially be achieved.
@@
    Classification Model Dependency: Enhancing the algorithm’s robustness by reducing its dependence on the effectiveness of classification model. A stronger classification model will leads to better results, particularly when the feature correlation is weak.
@@
    Avoiding Local Optima: Introducing mechanisms to prevent the algorithm from falling into local optima and ensuring that the model converges to an optimal solution.
These optimizations will further improve the robustness and applicability of the DCA, enabling it to handle a wider range of data characteristics and real-world scenarios, and provide even more reliable and accurate predictions.